OpenStudy (anonymous):

What is the y-intercept of the equation? y = 5(6.8)x 5 6.8 1

OpenStudy (shaik0124):

y=mx+c is the equation

OpenStudy (anonymous):

? :S

OpenStudy (anonymous):

if the prob would be like: y = 5(6.8)^x when x=0, y=y = 5(6.8)^0=5(1)=5 ans...

OpenStudy (anonymous):

also if the prob is like this y = 5^(6.8)x when x=0, y=y = 5^(6.8)(0)=5^0=1 ans..